OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = x +

OpenStudy (anonymous):

@abswag13

OpenStudy (anonymous):

@igr

OpenStudy (leader):

slope intercept form is y=mx+b

OpenStudy (leader):

So y=2x+b 4=2(3)+b 6+b=4 b=-2

OpenStudy (leader):

So it will be y=2x-2

OpenStudy (anonymous):

ok thnx can you help with 5 more?

OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y = x +

OpenStudy (anonymous):

4 i meant

OpenStudy (anonymous):

@Leader